4. Panther Tyres company Ltd. has announced initial public offering (IPO) of Rs. 50 per share. The company is not expected pay dividend but is expected to begin to do so in five years (at t = 5). The first dividend is expected to be Rs. 4.00 and to be received five years from today. That dividend is expected to grow at 6 percent into perpetuity. Suppose your required return is 10 percent. What is the estimated current intrinsic value? Will you subscribe for Panther Tyres IPO?


1
Expert's answer
2021-06-01T14:01:55-0400
"\frac{4}{50}\times100=8"

Since the first dividend will be 8%, and then there follows an increase in dividends by 6% annually ad infinitum, then, for example, over 20 years the average income will be


"\frac{1}{20}\times \frac{16+14\times6}{2}\times15=37.5"

This average income is well above the 10% target. I will subscribe to these promotions.